# Cost-sensitive computational adequacy of higher-order recursion in synthetic domain theory

### v1

@@ Appendix B Properties of predomains
Replete types are both complete and boundary separated because the latter can be defined by orthogonality conditions and is satisfied by $\Sigma$: it is complete by the axioms of SDT (Definition 3.16) and it is boundary separated by Proposition 3.15. Assuming Phoa’s principle, the proof that replete types are linked can be found in Taylor [31, Corollary 2.10] and Reus [22, Corollary 6.1.16]. Lastly, one can show that the objects whose link relation is antisymmetric can be defined as an orthogonality condition, thus we know that the link relation on every replete type is antisymmetric, hence it follows the intrinsic order on replete types are also partial orders since they are linked.
@@
Using the fact that $A$ is $\Sigma$-decidable, we obtain a map $f:(\P\vee A)\times(\P\vee A)\to\P\vee\Sigma$ such that $f(\eta_{\P\vee-}(x),\eta_{\P\vee-}(y))=\eta_{\P\vee-}(x=y)$ and $f(\star,-)=f(-,\star)=\star$. The desired characteristic map can then be defined as $\sigma\circ f$, where $\sigma:\P\vee\Sigma\to\Sigma$ is defined as follows:
@@ F.2 Fundamental lemma
Appendix G Model construction
See 9.1
Proof G.1.
Sterling and Harper [30] prove that every representable presheaf is a sheaf with respect to the finite cover topology, so we obtain an embedding $y:\mathscr{C}\hookrightarrow\mathrm{Sh}(\mathscr{C})$. Moreover, op. cit. shows that $\Sigma$ is a complete dominance in $\mathrm{Sh}(\mathscr{C})$. To show that $\mathrm{Sh}(\mathscr{C})$ also models the intension-extension phase distinction, we observe that the presheaf model of the intension-extension phase distinction of Niu et al. [17] restricts to a smaller model in $\mathscr{C}$: every subterminal object in $\widehat{\mathbb{I}}$ is an internal dcpo, and so we may take the subterminal $\P=\mathsf{y}_{\mathbb{I}}(\mathsf{ext})$ to be the phase proposition in $\mathscr{C}$. The phase proposition $\P=y(\P)$ in $\mathrm{Sh}(\mathscr{C})$ is classified by $\Sigma$: since $y$ is fully faithful, every $\phi:\Sigma$ arise from a unique map $1_{\widehat{\mathbb{I}}}\to\Omega_{\widehat{\mathbb{I}}}$.
Moreover, we can directly verify that $2_{\mathscr{C}}=2_{\widehat{\mathbb{I}}}=1_{\widehat{\mathbb{I}}}+1_{\widehat{1}}$ is internally orthogonal to $\P$ in $\mathscr{C}$. Under our chosen localization, $y$ preserves finite coproducts, so we have that $2_{\mathrm{Sh}(\mathscr{C})}=y(2_{\mathscr{C}})$. We observe that 2 is extensional because the restricted embedding $y:\mathscr{C}\to\mathrm{Sh}(\mathscr{C})$ is both full and faithful and preserves products. To see that $2$ is replete, we observe that 2 is isomorphic to its type of singletons:
Because $\Sigma$ is closed under finite joins, the type of singletons of 2 can be defined as the limit of a diagram of replete types, and so it is replete as well.

### v2

@@ Appendix B Properties of predomains
Replete types are both complete and boundary separated because the latter can be defined by orthogonality conditions and is satisfied by $\Sigma$: it is complete by the axioms of SDT (Definition 3.16) and it is boundary separated by Proposition 3.15. Assuming Phoa’s principle, the proof that replete types are linked can be found in Taylor [33, Corollary 2.10] and Reus [24, Corollary 6.1.16]. Lastly, one can show that the objects whose link relation is antisymmetric can be defined as an orthogonality condition, thus we know that the link relation on every replete type is antisymmetric, hence it follows the intrinsic order on replete types are also partial orders since they are linked.
@@
Using the fact that $A$ has $\Sigma$-equality, we obtain a map $f:(\P\vee A)\times(\P\vee A)\to\P\vee\Sigma$ such that $f(\eta_{\P\vee-}(x),\eta_{\P\vee-}(y))=\eta_{\P\vee-}(x=y)$ and $f(\star,-)=f(-,\star)=\star$. The desired characteristic map can then be defined as $\sigma\circ f$, where $\sigma:\P\vee\Sigma\to\Sigma$ is defined as follows: